# Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability

### v1

@@ 3.1 An alternative MLE-based strategy
Walker and Hjort, (2001) approached the problem of posterior consistency as follows. Taking $A_{\varepsilon}$ to be a Hellinger $\varepsilon$-ball around $f_{\theta_{\star}}$, consider
@@
The condition in (2) is often satisfied in regular parametric models, for instance as guaranteed by the classical theorem of Wilks, (1938). Nonetheless, this proof technique inherits the same core limitations as that of Walker, (1969): it relies on potentially stringent regularity conditions to guarantee MLE behavior (Kiefer and Wolfowitz,, 1956; van de Geer,, 2000; van der Vaart,, 2000), and achieves consistency through a sequence of technical bounds, rather than by appealing to a conceptual framework that directly explains why the posterior should concentrate. In the next section, we move beyond these approaches and develop a more principled and flexible route to consistency, rooted in the concept of sequential identifiability.
@@ 4 Sequential identifiability, oscillations and posterior consistency
As mentioned in the introduction, our starting point is the groundbreaking result of Schwartz, (1964). Given its central role, we state it here (without proof).
Theorem 1 (Schwartz,, 1964).
@@
While remarkable, Theorem 1 is not sufficient to address consistency in density estimation, as the topology induced by $d_{w}$ is too weak to meaningfully capture closeness between densities. In particular, it is possible to construct sequences of densities that do not converge in the Hellinger sense, yet whose associated cdfs converge pointwise to that of a continuous random variable—i.e., they converge weakly.
This phenomenon, which will be analyzed in greater detail in Theorem 3 and illustrated through a concrete model in Section 5, arises from the fact that a sequence of densities may oscillate indefinitely in a manner that precludes convergence in Hellinger distance, while still approximating the target cdf in distribution. In such cases, although the limit in density does not exist, the sequence may nonetheless approximate the cdf of the density around which it oscillates.
@@
Theorem 2 is noteworthy in that it establishes posterior consistency under very mild assumptions: sequential identifiability of the model and a prior that assigns positive mass to any KL neighborhood of the true parameter, the latter being a standard well-specification condition. The intuition behind the derivation of Theorem 2 is as follows: the KL support condition, via Schwartz’s theorem, ensures that the posterior concentrates in weak neighborhoods of $F_{\theta_{\star}}$. Sequential identifiability then rules out the possibility that such concentration occurs around points in the parameter space other than that corresponding to the true density, thereby yielding posterior consistency.
@@
A practical implication of this result is that, to establish parameter consistency at some $\theta_{\star}$, it suffices to verify that the model space—augmented by its weak limit points—does not contain $F_{\theta_{\star}}$ at more than one location (e.g., both within the parameter space and along a sequence $(\theta_{j})_{j\in\mathbb{N}}$ with $\|\theta_{j}\|\to\infty$). As we show in the next subsection, the possibility of such a scenario is tied to the presence of pathological oscillations of the model densities around $f_{\theta_{\star}}$, which is highly unlikely in any parametric modeling setting unless the modeler possesses specific knowledge of $f_{\theta_{\star}}$ and deliberately constructs the model to fail sequential identifiability at $\theta_{\star}$.
@@
Another insight from Theorem 2 arises from the following observations. If the posterior fails to be consistent at some $\theta_{\star}\in\Theta$, Theorem 2 implies the existence of a region in the augmented parameter space that induces sequential unidentifiability at $\theta_{\star}$, with the posterior accumulating with positive probability around such a weak limit point, separated from $\theta_{\star}$ in the Euclidean sense. Under the basic assumption of identifiability, however, inconsistency cannot result in posterior concentration around a different point within the parameter space, as this would contradict identifiability. Rather, the posterior mass must shift toward a region in the augmented space, lying outside the original parameter space, where the true distribution is still recovered. In other words, the posterior does not simply “miss” the true parameter by concentrating around a nearby but incorrect value; instead, it shifts toward regions in the augmented space proper, such as points at infinity, which correspond to weak limits of the true distribution. This observation suggests a practical heuristic for diagnosing strong consistency: in addition to verifying sequential identifiability, one may examine whether the posterior remains confined within the parameter space, rather than drifting toward, for instance, infinity. In light of our discussion, such stability would provide evidence in support of consistency.
@@ 4.1 The role of oscillations
While sequential identifiability is the central concept of our theoretical analysis, it is crucial to understand the implications of its failure. The next theorem addresses this question by examining the behavior of a sequence of density functions that do not converge in the Hellinger metric (which, in our parametric setting, is assumed to be equivalent to non-convergence of the associated parameter sequence), yet whose corresponding distribution functions do converge to a proper limit.
@@
The preceding result implies that, if sequential identifiability fails at some $\theta\in\Theta$, then the model must contain a sequence of densities that oscillate arbitrarily frequently around $f_{\theta}$. Figure 1 provides a graphical illustration of how such a sequence would behave. Consequently, ruling out this type of pathological behavior is sufficient to guarantee posterior consistency. In practice, posterior inconsistency due to such behavior could only arise if the modeler possessed precise knowledge of the true density and deliberately introduced carefully designed oscillations around it. We observe that this is an implausible scenario in any reasonable parametric modeling context, which makes our conditions for parametric consistency virtually universal.
@@
Figure 1: The initial segment of a sequence $f_{1},f_{2},\dots$ of indefinitely oscillating densities (solid gray, top row), whose corresponding sequence of cdfs (solid gray, bottom row) converges to a proper limiting distribution (dashed black). Such convergence is only possible because the density sequence exhibits increasingly frequent oscillations around the density associated with the limiting cdf.
@@ 5 An illustrative model
$F_{\theta}(x)=\frac{x+\sin(\theta x)/\theta}{1+\sin(\theta)/\theta}.$
@@
    For any $\theta\geq 0$ and prior $\Pi$ on $\Theta$, if $\Pi(A)>0$ for every neighborhood $A$ of $\theta$, then $\theta\in\textnormal{KLS}(\Pi)$;
@@
The first property confirms that the assumption of equivalence between the Euclidean and Hellinger metrics, made throughout this article, holds in this setting as well. This allows us to move freely between the two notions of convergence without loss of generality. The second property implies that, for the KL support condition to hold at all parameter values, it suffices for the prior to have full support on $\Theta$—a very mild requirement. Combined with the third property, this enables us to invoke Theorem 2 and obtain the following corollary.
@@
This result is remarkable because, as we demonstrate in Subsection 5.1 below, the family $\mathcal{F}_{\Theta}$ fails to satisfy classical sufficient conditions for consistency proposed in earlier literature (cf. Section 3). Nonetheless, our simple criterion of sequential identifiability seamlessly yields posterior consistency at all $\theta>0$.
Due to the fourth property in Proposition 1, however, the same argument does not extend to $\theta=0$, as the model is sequentially unidentifiable at that point. In particular, while a prior with full support guarantees posterior concentration in weak neighborhoods of $F_{0}$ (thanks to Schwartz’s theorem), such concentration may occur in the “wrong” region of the parameter space—namely, toward infinity. In line with Theorem 3, this happens because this cosine-based model oscillates around the uniform density with increasing frequency as the parameter $\theta$ diverges. It is worth emphasizing the specificity of the model’s construction in producing such behavior: the oscillations are finely tuned to occur around the uniform density and do not induce inconsistency at any other parameter value. This further reinforces our earlier argument that, although violations of sequential identifiability are theoretically possible, they are highly unlikely to arise in any reasonably specified parametric model, let alone pose a threat to posterior consistency. Indeed, sequential unidentifiability at the true parameter value would require the modeler to possess precise knowledge of the data-generating process and to precisely engineer the statistical model to behave pathologically around it—as we have done in this example.
@@ 5.1 Failure of classical conditions
More fundamentally, we next establish that, under this model, the MLE is inconsistent at all data-generating parameter values. This result reinforces the limitations of the approach developed by Walker, (1969), which relies on model regularity to infer posterior consistency from MLE consistency. The inconsistency of the MLE renders the methodology of Walker and Hjort, (2001) inapplicable as well, since that framework also presupposes MLE consistency.
@@ 5.2 Consistency at the uniform density
The first approach we consider is the one introduced in the seminal papers by Ghosal et al., (1999) and Barron et al., (1999), which establishes exponential decay of the posterior mass outside Hellinger neighborhoods of $\theta_{\star}$ by splitting the complement of such neighborhoods into two $n$-dependent regions: the first has slowly growing complexity—e.g., measured in terms of Hellinger metric or upper-bracketing entropy (van de Geer,, 2000; Wainwright,, 2019)—forcing the likelihood ratio to decay exponentially; the second region is directly assumed to have exponentially small prior (hence posterior) mass. Together with the KL support assumption and an upper-bound of $d_{h}(f_{\theta},f_{\theta^{\prime}})$ by the corresponding Euclidean distance (see the supplementary material for more details), this leads to the following result.
@@ 6 Discussion
In this work, we introduced a general framework for studying posterior consistency in parametric models, based on the simple yet powerful notion of sequential identifiability. Departing from the regularity-based assumptions commonly found in the literature, our approach builds on the observation that, by the seminal result of Schwartz, (1964), weak consistency is guaranteed under a mild KL support condition. Sequential identifiability then emerges as a minimal additional assumption that enables the translation of weak consistency into parameter consistency. We further showed that inconsistency can arise only in the presence of oscillatory behavior in the model—an insight that, together with the sequential identifiability condition itself, leads to consistency results under assumptions significantly weaker than those required by classical theory. This perspective enables the study of models beyond the reach of existing approaches, including the parametric family we constructed and analyzed in detail.
@@
Our work has a multiplicity of ramifications. Most directly, it offers novel and accessible guidance for the applied Bayesian statistician seeking to understand the asymptotic behavior of parametric models of interest. Building on our analysis of sequential identifiability, our message to the modeler is straightforward: rather than verifying a list of regularity conditions, simply avoid introducing arbitrarily oscillating densities into the model in the first place. If, for some reason, such oscillations are unavoidable, there is still little cause for concern about inconsistency—unless one possesses precise knowledge of the true data-generating distribution and is willing to meticulously design pathological oscillations centered around it. However, in such a scenario the need for statistical inference itself becomes questionable: after all, if the modeler already knows so much about the true distribution as to craft these adversarial features, inference would hardly be necessary. As we have illustrated, our carefully constructed example required exactly this kind of contrived design to challenge consistency, yet it remained sequentially identifiable at all but a single parameter value. Moreover, even at the parameter value where sequential identifiability fails, we showed that posterior consistency can still be salvaged through a fine-grained analysis of the model’s oscillatory behavior and suitable control via the prior.
@@
Second, our analysis bears strong connections with the literature on nonparametric posterior consistency. Although not explicitly, the sieve complexity conditions introduced by Barron et al., (1999), Ghosal et al., (1999), and Walker, (2004) aim to address a similar underlying issue: in those works, oscillatory behavior in the density model—leading to sequential unidentifiability—is quantified and controlled, for instance, via the Hellinger metric entropy of sets where most of the prior mass is concentrated. However, it is worth emphasizing that, in nonparametric models, sequential unidentifiability is almost always inherent. For example, infinite Gaussian mixtures can weakly approximate any distribution on the real line while failing to converge in density, due to components with rapidly increasing precision parameters. In such settings, sequential identifiability is not a reasonable assumption, and a more refined analysis of the interaction between the model’s oscillatory behavior and the prior is necessary.
@@
Finally, our analysis—particularly the construction of a model based on cosine oscillations—sheds new light on a well-known counterexample by Barron et al., (1999), in which the posterior is weakly consistent but not consistent in Hellinger distance. In that example, positive prior mass is placed on each member of a family of discontinuous, oscillatory densities on $[0,1]$ that alternate between values of 0 and 2 across disjoint intervals, thereby weakly approximating the uniform density. In our terminology, this corresponds to a model that is sequentially unidentifiable at the uniform density, and the mechanism driving inconsistency is closely related to our example, in which continuous oscillations occur between 0 and (approximately) 2 around the uniform density. Nevertheless, our model has been shown to achieve consistency at the uniform distribution under mild tail conditions on the prior, while the carefully placed prior mass on discontinuous, oscillatory densities in the Barron et al., (1999) example results in inconsistency. Although beyond the scope of this study, these observations suggest the possibility that additional mild assumptions on the nature of oscillations (e.g., continuity) or on the prior (e.g., no positive mass on single parameters or densities) may help to rule out inconsistent posterior behavior even in nonparametric settings.
@@
Supplementary Material for “Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability"
@@ Appendix D Proof of Proposition 2
$\int_{0}^{1}f_{\theta_{\star}}(x)\ln\left(\hbox{$\frac{1}{2}$}f_{\theta_{\star}}(x)\right)\mathrm{d}x$ $\leq\int_{0}^{1}f_{\theta_{\star}}(x)\left(\hbox{$\frac{1}{2}$}f_{\theta_{\star}}(x)-1\right)\mathrm{d}x$
@@ Appendix H Proof of Theorem 5
For the second factor in Equation (H), one obtains

### v2

@@ 3.1 An alternative MLE-based strategy
Alternatively, Walker and Hjort, (2001) approached the problem of posterior consistency as follows. Taking $A_{\varepsilon}$ to be a Hellinger $\varepsilon$-ball around $f_{\theta_{\star}}$, consider
@@
The condition in (2) is often satisfied by regular parametric models, for instance as guaranteed by the classical theorem of Wilks, (1938). Nonetheless, this proof technique inherits the same core limitations as that of Walker, (1969): it relies on potentially stringent regularity conditions to guarantee MLE behavior (Kiefer and Wolfowitz,, 1956; van de Geer,, 2000; van der Vaart,, 2000), and achieves consistency through a sequence of technical bounds, rather than by appealing to a conceptual framework that directly explains why the posterior should concentrate. In the next section, we move beyond these approaches and develop a more principled and flexible route to consistency, rooted in the concept of sequential identifiability.
@@ 4 Sequential identifiability, oscillations and posterior consistency
As mentioned in Section 1, our starting point is the groundbreaking result of Schwartz, (1964). Given its central role, we formally state it here (without proof).
Theorem 1 (Schwartz).
@@
While remarkable, Theorem 1 is not sufficient to address consistency in density estimation, as the topology induced by $d_{w}$ is too weak to meaningfully capture closeness between densities. In particular, it is possible to construct sequences of densities that do not converge in the Hellinger sense, yet whose associated CDFs converge pointwise to that of a continuous random variable—i.e., they converge weakly. This phenomenon, which will be analyzed in greater detail in Theorem 3 and illustrated through a concrete model in Section 5, arises from the fact that a sequence of densities may oscillate indefinitely in a manner that precludes convergence in Hellinger distance (hence in terms of parameter values), while still approximating the target CDF in distribution (see Figure 1 in Subsection 1.1). In such cases, although the limit in density does not exist, the sequence may nonetheless approximate the CDF of the density around which it oscillates.
@@
It is worth noting that this seemingly “obvious” extension of the notion of identifiability has been largely overlooked for a simple reason: it offers no benefit for analyzing the convergence properties of the MLE. Because, as we have demonstrated, Bayesian consistency has traditionally been studied in close connection with MLE consistency, our notion in Definition 3 would not have surfaced within that framework. However, as we show next, if sequential identifiability holds, it leads directly to Bayesian consistency. This observation highlights a key message of our work: Bayesian consistency and MLE consistency should be treated using fundamentally different tools, contrary to the prevailing approach in the literature. This point will be further illustrated in Section 5 with a concrete parametric model.
@@
Theorem 2 is noteworthy in that it establishes posterior consistency under quite mild conditions: sequential identifiability of the model and a prior that assigns positive mass to any KL neighborhood of the true parameter, the latter being a standard well-specification assumption. As we illustrate in Section 5, this condition captures the core mechanism behind Bayesian consistency and, unlike the classical assumptions in the literature, is entirely decoupled from MLE convergence. In fact, our illustrative model will reveal that even when sequential identifiability holds at a given parameter value, ensuring posterior consistency at it, the MLE may still be inconsistent due to likelihood peaks at the data. The intuition behind the derivation of Theorem 2 is as follows: the KL support condition, via Schwartz’s theorem, ensures that the posterior concentrates in weak neighborhoods of $F_{\theta_{\star}}$. Sequential identifiability then rules out the possibility that such concentration occurs around points in the parameter space other than that corresponding to the true density, thereby yielding posterior consistency.
@@
A practical implication of this result is that, to establish parameter consistency at some $\theta_{\star}$, it suffices to verify that the model space—augmented by its weak limit points—does not contain $F_{\theta_{\star}}$ at more than one location (e.g., both within the parameter space and along a sequence $(\theta_{j})_{j\in\mathbb{N}}$ with $\lim_{j\to\mathbb{N}}\|\theta_{j}\|=\infty$). As we show in the next subsection, the possibility of such a scenario is tied to the presence of pathological oscillations of the model densities around $f_{\theta_{\star}}$, which is highly unlikely in any parametric modeling setting unless the modeler possesses specific knowledge of $f_{\theta_{\star}}$ and deliberately constructs the model to fail sequential identifiability at $\theta_{\star}$. Notice also that, in line with our earlier discussion in Subsection 1.1, a brief inspection of the proof of Theorem 2 reveals that its validity is not restricted to the case of a finite-dimensional Euclidean parameter space $\Theta$, and in fact formally extends to nonparametric models as well. Nevertheless, as we have argued, sequential identifiability is generally not a tenable assumption in nonparametric settings, where, unlike in the parametric case, sequential unidentifiability is typically inherent.
@@
Another insight from Theorem 2 arises from the following observations. If the posterior fails to be consistent at some $\theta_{\star}\in\Theta$, Theorem 2 implies the existence of a region in the augmented parameter space that induces a lack of sequential identifiability at $\theta_{\star}$, with the posterior accumulating with positive probability around such a weak limit point, separated from $\theta_{\star}$ in the Euclidean sense. Under the basic assumption of (traditional) identifiability, however, inconsistency cannot result in posterior concentration around a different point within the parameter space, as this, together with Schwartz’s assurance of weak consistency, would contradict identifiability. Rather, the posterior mass must shift toward a region in the augmented space, lying outside the original parameter space, where the true distribution is also recovered. In other words, the posterior does not simply “miss” the true parameter by concentrating around a nearby but incorrect value; instead, it shifts toward regions in the augmented space proper, such as points at infinity, which correspond to weak limits of the true distribution. This observation suggests a practical heuristic for diagnosing strong consistency: in addition to directly verifying sequential identifiability, one may examine whether the posterior remains confined within reasonable regions of the parameter space, rather than drifting toward, for instance, infinity. In light of our discussion, such stability would provide evidence in support of consistency.
@@ 4.1 The role of oscillations
While sequential identifiability is the central concept of our theoretical analysis, it is crucial to understand the implications of its failure. The next theorem addresses this question by examining the behavior of a sequence of density functions that do not converge in the Hellinger metric (which, in our parametric setting, is equivalent to non-convergence of the associated parameter sequence), yet whose corresponding distribution functions do converge to a proper limit.
@@
The preceding result implies that, if sequential identifiability fails at some $\theta\in\Theta$, then the model must contain a sequence of densities that oscillate arbitrarily frequently around $f_{\theta}$. Figure 1 in Subsection 1.1 provides a graphical illustration of the behavior of such a sequence. Consequently, ruling out this kind of pathological behavior is sufficient to guarantee posterior consistency. In practice, due to the inherently limited expressive power of finite-dimensional models (with respect to weakly approximating distributions outside the proper parameter space), posterior inconsistency from this mechanism could arise only if the modeler had precise knowledge of the true density and intentionally introduced carefully constructed oscillations around it. Notice that these oscillations would not only need to be present, but also to integrate in such a way as to yield the correct distribution function away from the true parameter value. Clearly, such a contrived construction is implausible in any real-world parametric modeling scenario, rendering our simple sequential identifiability condition effectively universal for parametric consistency.
@@ 5 An illustrative model
$F_{\theta}(x)=\frac{x+\sin(\theta x)/\theta}{1+\sin(\theta)/\theta},\quad\forall x\in[0,1].$
@@
    For any $\theta\geq 0$ and prior $\Pi$ on $\Theta$, if $\Pi(A)>0$ for every Euclidean neighborhood $A$ of $\theta$, then $\theta\in\textnormal{KLS}(\Pi)$;
@@
The first property confirms that the assumption of equivalence between the Euclidean and Hellinger metrics, made throughout this article, holds in this setting as well. This allows us to move freely between the two notions of convergence without loss of generality. The second property implies that, for the KL support condition to hold at all parameter values, it suffices for the prior to have full support on $\Theta$. Combined with the third property, this enables us to invoke Theorem 2 and obtain the following corollary.
@@
This result is remarkable because, as we demonstrate in Subsection 5.2 below, the family $\mathcal{F}_{\Theta}$ fails to satisfy the classical sufficient conditions for consistency proposed in earlier literature (cf. Section 3). Nonetheless, our simple criterion of sequential identifiability seamlessly yields posterior consistency at all $\theta>0$.
Due to the fourth property in Proposition 1, however, the same argument does not extend to $\theta=0$, as the model is sequentially unidentifiable at that point. In particular, while a prior with full support guarantees posterior concentration in weak neighborhoods of $F_{0}$ (thanks to Schwartz’s theorem), such concentration may occur in the “wrong” region of the parameter space—namely, toward infinity. In line with Theorem 3, this happens because the cosine-based model oscillates around the uniform density with increasing frequency as the parameter $\theta$ diverges.
5.1 Some remarks on the illustrative model
Before continuing the formal analysis of our example, a few important remarks about its construction are in order. First, the density model defined by (3) demonstrates that, although sequential unidentifiability is theoretically possible, it is extremely unlikely to occur in any reasonably specified parametric model, and thus poses little practical threat to posterior consistency. In our example, inconsistency can arise only if the true parameter is $\theta_{\star}=0$, a scenario we engineered by precisely tailoring cosine-based oscillations around the corresponding (uniform) density. Such a construction is implausible in real-world settings: not only are densities with pathologically frequent oscillations rarely found in practical parametric models, but even if present, it is highly unlikely that they would align so precisely with the true (and unknown) density. Since a parametric model has only limited capacity to generate these oscillations around any given target, the conditions for inconsistency require both unlikely model structure and foreknowledge of the data-generating process—making sequential unidentifiability an essentially self-inflicted phenomenon for the purposes of parametric inference.
A more subtle point arises by noticing that our construction can be generalized by creating problematic oscillatory behavior around an entire parametric family, rather than just the uniform distribution used in our example. For instance, [^9] The following construction starts from a simple parametric family $\mathcal{G}$ to keep a clear notation, though it can be further extended to more general families. consider the parametric family $\mathcal{G}$ containing all densities of the form
$g_{\lambda}(x)=\lambda^{-1}1_{[0,\lambda]}(x)$
for $\lambda\in\Lambda\subseteq[1,2]$, and construct a new model $\mathcal{H}$ containing all densities of the form
[…]
for $\lambda\in\Lambda$ and $\omega\in\Omega:=[0,\infty)$, where $\mu\in(0,1)$ is a constant small enough (depending on $\Lambda$) to ensure positivity of all members of $\mathcal{H}$. Our illustrative example is then recovered if $\Lambda=\{1\}$ and $\mu=1$. The associated CDFs satisfy
$H_{\lambda,\omega}(x)=\frac{x/\lambda+\mu\sin(\omega x)/\omega}{1+\mu\sin(\omega\lambda)/\omega}$
[…]
@@ 5.2 Failure of classical conditions
More fundamentally, we next establish that, under this model, the MLE is inconsistent at all data-generating parameter values. This result reinforces the limitations of the approach developed by Walker, (1969), which relies on model regularity to infer posterior consistency from MLE consistency. The inconsistency of the MLE renders the methodology of Walker and Hjort, (2001) inapplicable as well, since that framework also presupposes regular MLE behavior.
@@ 5.3 Consistency at the uniform density
The first approach we consider is the one introduced in the seminal papers by Ghosal et al., (1999) and Barron et al., (1999), which establishes exponential decay of the posterior mass outside Hellinger neighborhoods of $\theta_{\star}$ by splitting the complement of such neighborhoods into two $n$-dependent regions: the first has slowly growing complexity—e.g., measured in terms of Hellinger metric or upper-bracketing entropy (van de Geer,, 2000; Wainwright,, 2019)—forcing the likelihood ratio to decay exponentially; the second region is directly assumed to have exponentially small prior (hence posterior) mass. Together with the KL support assumption and an upper-bound on $d_{h}(f_{\theta},f_{\theta^{\prime}})$ by the corresponding Euclidean distance (see the supplementary material for more details), this leads to the following result.
@@ 6 Discussion
We introduced a general framework for studying posterior consistency in parametric models, centered on the simple yet powerful notion of sequential identifiability. Departing from the classical approach that links posterior consistency to the regular behavior of maximum likelihood estimation, our perspective builds on the foundational result of Schwartz, (1964), which guarantees weak consistency under a mild Kullback-Leibler support condition. Sequential identifiability then arises as the minimal additional assumption needed to lift weak consistency to consistency at the level of parameters. We also showed that inconsistency can only occur when the model admits self-inflicted pathological oscillatory behavior around the true density. This insight, along with the identifiability criterion itself, allows us to establish consistency under assumptions significantly less restrictive than those required by classical regularity-based theories, which were primarily designed to prevent MLE overfitting as a consequence of likelihood peaks at the data points. Our framework thus opens the door to analyzing models that fall outside the scope of existing results, including the parametric example we constructed and examined in detail.
@@
Our work has a multiplicity of ramifications. Most directly, it offers novel and accessible guidance for the applied Bayesian statistician seeking to understand the asymptotic behavior of parametric models of interest. Building on our analysis of sequential identifiability, our message to the modeler is straightforward: rather than verifying a list of regularity conditions, simply avoid introducing arbitrarily oscillating densities into the model in the first place. Even when such oscillations are unavoidable, parametric inconsistency remains unlikely unless the modeler is able to design oscillations that align with the unknown data-generating process. However, in that scenario, statistical inference under a parametric model becomes questionable. Either the modeler is confident in using a finite-dimensional family and yet somehow knows the true distribution well enough to adversarially target it with pathological oscillations, making inference hardly necessary; or the modeling goal demands such flexibility that sequential unidentifiability arises over a broad subset of the parameter space, in which case a nonparametric approach may be more suitable. Our illustrative example required exactly this kind of contrived construction to potentially induce inconsistency, and even then, only at a single, sequentially unidentifiable parameter value. Remarkably, we further showed that consistency can still be recovered at that value through a detailed analysis of the model’s oscillatory behavior and appropriate control via the prior. This was done by means of techniques inspired by the study of nonparametric consistency, where this kind of extreme oscillatory behavior naturally belongs.
@@
Second, as we have discussed in detail, our analysis bears strong connections with the literature on nonparametric posterior consistency. Specifically, by recognizing some fundamental differences between parametric and nonparametric models, we have arrived at sequential identifiability as a key condition for parametric consistency, while recognizing its inappropriateness in nonparametric settings. Nevertheless, our analysis on oscillations, motivated by the need to understand the implications of the failure of sequential identifiability, is intimately tied with the most common approaches to nonparametric consistency. Although not explicitly, the sieve complexity conditions introduced by Barron et al., (1999), Ghosal et al., (1999), and Walker, (2004) aim to address the same underlying issue: in those works, oscillatory behavior in the density model is quantified and controlled, for instance, via the Hellinger metric entropy of sets where most of the prior mass is concentrated. This is analogous to our calculations in Section 5, where we established posterior consistency at the only parameter value for which our illustrative model is not sequential identifiable.
@@
Finally, our analysis—particularly the construction of a model based on cosine oscillations—sheds new light on a well-known counterexample by Barron et al., (1999), in which the posterior is weakly consistent but not consistent in Hellinger distance. In that example, positive prior mass is placed on each member of a family of discontinuous, oscillatory densities on $[0,1]$ that alternate between values of 0 and 2 across disjoint intervals, thereby weakly approximating the uniform density. In our terminology, this corresponds to a model that is sequentially unidentifiable at the uniform density, and the mechanism driving inconsistency is closely related to our example, in which continuous oscillations occur between 0 and (approximately) 2 around the uniform density. Nevertheless, our model has been shown to achieve consistency at the uniform distribution under mild tail conditions on the prior, while the carefully placed prior mass on discontinuous, oscillatory densities in the example of Barron et al., (1999) results in inconsistency. Although beyond the scope of this study, these observations suggest the possibility that additional mild assumptions on the nature of oscillations (e.g., continuity) or on the prior (e.g., no positive mass on single parameters or densities) may help to rule out inconsistent posterior behavior even in nonparametric settings.
To summarize our findings, for a parametric model to be inconsistent, not only must a distribution $F_{\infty}$ exist at the weak boundary of the model—say, as the parameter diverges to infinity—but this distribution must also coincide with the one from which the sample is generated. As we have shown, this alignment can only result from an implausible and adversarial model design involving oscillations carefully tailored around the true, unknown density. Even in such cases, posterior inconsistency does not arise unless the prior assigns sufficient mass to these oscillatory features. For instance, in a finite Gaussian mixture model with $K$ components, the boundary distributions $F_{\infty}$ corresponds to discrete measures with at most $K$ atoms, which can be ruled out as plausible true distributions because they do not admit a density. In our cosine-based example, where the oscillations are deliberately engineered so that $F_{0}=F_{\infty}$ correspond to the uniform distribution, we showed that inconsistency at $\theta=0$ is still ruled out even under priors with heavier-than-Cauchy tail, due to insufficient mass being placed on the problematic region. In short, parametric models, unlike nonparametric ones, may be regarded as universally consistent unless one adopts a highly artificial construction that simultaneously entangles the true distribution, the density model, and the prior.
@@
Supplementary Material for “Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability”
@@
$\int_{0}^{1}f_{\theta_{\star}}(x)\ln\left(\frac{1}{2}f_{\theta_{\star}}(x)\right)\mathrm{d}x$ $\leq\int_{0}^{1}f_{\theta_{\star}}(x)\left(\frac{1}{2}f_{\theta_{\star}}(x)-1\right)\mathrm{d}x$
@@
For the second factor in Equation (6), one obtains